OpenStudy (anonymous):

On a trip, the fare was 50¢ for each adult and 25¢ for each child. If 30 passengers paid $12.25, how many adults and children went?

OpenStudy (jhannybean):

Another system of linear equations.

OpenStudy (jhannybean):

adult = x, child = y total = 30 so x + y = 30 0.50x + 0.25y = 12.25 use elimination method (preferably) to solve

sammixboo (sammixboo):

Yes, you explained it a lot easier than me, ahaha :)

OpenStudy (kamibug):

12.25 = 0.50a + 0.25c a + c = 30 a = 30 - c 12.25 = 0.50(30-c) + 0.25c 12.25 = 15 - 0.50c + 0.25c 12.25 = 15 - 0.25c -2.75 = -0.25c c = 11 11 children went. Now to solve for adults just do 30 - 11 = ?